A simple pendulum has a time period of 1.42 s on Earth. The gravitational field strength at the surface of Mars is 0.37 times that at the surface of the Earth. What is the time period of the pendulum on Mars?


1
Expert's answer
2021-02-24T12:51:50-0500
"T_M=2\pi\sqrt{\dfrac{L}{0.37g}}=\dfrac{1}{0.608}\cdot2\pi\sqrt{\dfrac{L}{g}}=\dfrac{T_E}{0.608}.""T_M=\dfrac{1.42\ s}{0.608}=2.33\ s."